solve for u (7+u)^2 =2u^2+18u+52
Answer #1

( 7 + u )^2 = 2u^2 + 18u + 52

expanding left hand side

49 + 14u + u^2 = 2u^2 + 18u + 52

bringing like terms on one side

u^2 + 4u + 3 = 0

( u + 3) ( u +1 ) = 0

u = - 3 , -1